For the space truss shown below, there is a Ball-and-socket joint at A. At C, only support reaction is in z direction and at B, there are support reactions in y and z directions. A force P=10k kN is applied to joint D. Triangle ABC is in x-y plane at z=0. For this truss,

(a) Determine the sequence of joints with which the entire truss can be solved for the member forces and support reactions using the method of joints.

(b) Solve the first joint only for forces in members connected to the joint.

Note: For part (a), only the correct sequence of joints is needed as an answer (e.g. A,B,C,D). You do not need to solve the entire truss.

N+ D P 8m с A 1m 5m x 7m B 10m
Answer #1

Solution g) Draw FBD of truss and define the sequence to solve using method of joints:- Note:a. using method of joint we can

i BLD i fclo= -så-Ĵ - BR 2 Alp = 0.53ê -0.11) -0.84 sî +69 - 8 118) +6397682 suj -0.72 res)²+(-1)2 +682 = -0.53î - allŷ -0.BuThanks